13 November 2007

RHIC Startup:

Working on blue ramp.  Try to do ramp development with tune feedback.  Need to have Larry there to do this work.  Angelika will correct the orbit in Blue, for however long this store lasts. (Oops, store was accidentally tossed.)  At the end of this store, Controls will go in to replace the DAC board.  There is a water leak in the 912 cooling tower for Booster RF.  There will be a 2 hour down time during which the cooling water leak will be repaired, and the DAC board will be replaced with a spare.

Yellow PS work, waiting for ramp to be fixed.  Steve will make the yellow ramp at slow factor of 1 for 90% of the ramp, and slow factor 2 for the last 10%.  Steve will create a new ramp, dAu82, that extends the beta squeeze on Yellow.  Blue ramp will be unchanged  Steve estimates that this work will take ~1 day.  But we still have large radial excursions, as we go up the ramp.  It is better to fix the ramp than to put in trim magnet settings which would be very big.  So Steve will modify the ramp instead.  The Yellow ramp that is currently installed has an 8 mm orbit excursion at transition.   

RF needs time for conditioning the cavities in RHIC.  Total conditioning time required is 48 hours, with a minimum block of time being 2 hours.  Yellow can be conditioned when beam is in Blue, and vice versa.  This needs to be coordinated with all other beam development work.

F3 problem last week was due to oxidation of connector.  It was cleaned and replaced and it works fine.  F6 also has problems.  Spare parts are ordered.  Pulsed Power Group is down 3 FTEs last week, and this will continue to be a problem until after Thanksgiving.

Blue Main Quad PS is off by 0.5 Amps. This may be due to a quench last week.

A Flow switch turned off the MMPS when it went out of adjustment.  It has been readjusted.

 



Scheduling Physicist: Mike Sivertz Last updated:
Email:sivertz@bnl.gov
Office Phone: 631-344-6102
Digital Pager: 7528 or 4300
Home Phone: 631-878-9020